OpenStudy (anonymous):

Solve using the addition method. What is the x-value of the solution to the system? 2x – 3y = -1 3x + 4y = 24

OpenStudy (amistre64):

we can multiply an equation and it keeps its inherent properties intact

OpenStudy (amistre64):

to get rid of the ys, we need to multiply the top and bottom by different values; something such that they will equal the same thing (but opposite)

OpenStudy (amistre64):

or we can go fractiony :)

OpenStudy (amistre64):

2x – 3y = -1 ; /3 3x + 4y = 24 ; /4 2x/3 – 3y/3 = -1/3 3x/4 + 4y/4 = 24/4 2x/3 – y = -1/3 3x/4 + y = 6 --------------- (2/3 + 3/4)x = 6 - 1/3 \[x=\frac{6-\frac{1}{3}}{\frac23+\frac34}\]

OpenStudy (anonymous):

I have no idea what you just did

OpenStudy (amistre64):

lol, um, i divided the top by 3 to make a -y i divided the bottom by 4 to make a +y when you add the 2 resulting equations together, the ys cancel; and you solve for x

OpenStudy (amistre64):

we can do this the same way, but multiply the top by 4 and the bottom by 3

OpenStudy (amistre64):

2x – 3y = -1 ; *4 3x + 4y = 24 ; *3 8x – 12y = -4 9x + 12y = 72 -------------- 17x = 68

OpenStudy (amistre64):

the addition method means that you want to eliminate one of the variables; to do that, you have to multiply (or divide) by some value that turn one of them into opposites

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

I understand that one now. Could you help me with another?

OpenStudy (amistre64):

post it up on the left, and ill see if i can get around to it

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (amistre64):

youre welcome :)